kabaribeau:

Simplify the ratio 4lb over 10 oz. Use the conversion 16 oz. = 1 lb

Shadow:

Hello

Shadow:

The most you can reduce this ratio by is by a factor of 2, which they both share. \[\frac{ 4 }{ 10 } \div \frac{ 2 }{ 2 } = \frac{ 2lb }{ 5oz }\]

Vocaloid:

I believe they would want you to use the conversion factor to convert 4 lbs to oz. and then re-write the ratio in simplest form